there are 9 women and 7 men in our class. how many different 4 person committees are possible if we want an equal number of men and women on the committees?

Answer by tommyt3rd(5050)

Ways to choose a woman: 9*8=72
Ways to choose a man: 7*6=42
Ways to choose: 72*42=3024

:)